2. Find the grouped means for the ages of male and females listed belovw Male 12 10 Female Age <15 years 15-24 years 25-54 years 55-64 years 65+ years 10 31 6 3. Find the unweighted and weighted mean centers for the following settlements Settlement (x, y) Population 1,000 50,000 2,500 3,000 8,000 (7.8, 2.3) (9.4, 3.3) (8.8, 3.7)
Answer #1

2) solution :

Given data:

Ages Male Female
< 15 years 12 10
15-24 years 10 9
24-54 years 33 31
55-64 years 8 8
65+years 6 8

Now, we have to fin dout the:

Grouped means for both male and females :

We already know the formula for the  Grouped mean is,

  

Here m = mid point

f = frequency

n = total

Now, Grouped means for males:

Ages Male (f) mid point ( m) mf
< 15 years 12 15/2 = 7.5 7.5 *12= 90
15-24 years 10 (15+24) /2 = 19.5 19.5*10= 195
25-54 years 33 ( 25+54 )/2 =39.5 39.5*33 = 1303.5
55-64 years 8 (55+64)/2 = 59.5 59.5*8 = 476
65+years 6 (65+100)/2 = 82.5 82.5*6 = 495
Total 69 mf = 2559.5

Now, Grouped mean   

= 2559.5 / 69

=37.094

Grouped mean for males  = 37.094

Now, we fin dout the:

Grouped mean for Females :

Ages Female (f) mid point ( m) mf
< 15 years 10 15/2 = 7.5 7.5*10= 75
15-24 years 9 (15+24) /2 = 19.5 19.5*9 =175.5
24-54 years 31 ( 25+54 )/2 =39.5 39.5*31= 1224.5
55-64 years 8 (55+64)/2 = 59.5 59.5* 8 = 476
65+years 8 (65+100)/2 = 82.5 82.5*8 = 660
Total 66 mf = 2611

Now, Grouped mean

= 2611 / 66

= 39.5606

Grouped mean for Females  =39.5606

Note : As per the HOMEWORKLIB RULES, one question is enough, so i am answered only one question if you want remaining question please re upload as another question.

Thank you,